OpenStudy (anonymous):

A 2.300g sample of CuxCly*zH2O weighed only 1.950g after heating.After aluminumwas added,a total of 1.234g of solid copper was collected. what is the formula of the hydrate?

OpenStudy (jebonna):

First work out the grams of each molecule in the substance. Total g of the sample: 2.300g Copper weight: 1.234g H2O weight: 2.300-1.950 = 0.350g (using 1.950 because that is the weight after heating, meaning the water would have dissolved) Chlorine weight: 1.950 - 1.234 = 0.716g (using 1.950 because that is the weight without water, meaning it is just Cu and Cl left. I am also using 1.234 because that is the weight of JUST the copper. This means taking one away from the other would equal chlorine) Now work out the empirical formula. Work out moles of Cu and Cl Moles = mass/molar mass Cu = 1.234/63.5 = 0.0194 (3 sig.fig) Cl = 0.716/35.5 = 0.0202 (3 sig.fig) Divide smallest moles by each answer. Cu = 0.0194/0.0194 = 1 Cl = 0.0202/0.0194 = 1 (rounded down) Therefore the formula is CuCl Now do the empirical formula including H2O Work out moles of H2O and CuCl and divide the smallest by each other and you will get how many H2O's are in the hydrate. (You will have to work this out, I have to leave D: ) I think that is right, but get a second opinion just in case.